OpenStudy (anonymous):

Which of the following is not a step used when graphing the equation 2x + y = 14 using the x- and y- intercepts method? Count down 2 and right 1 from the y-intercept. Plot the y-intercept (0, 14). Plot the intercept (7, 0). Draw a line through the two points.

OpenStudy (anonymous):

>:/ my answer didn't post here we go again...

OpenStudy (anonymous):

The equation 2x+y=14 can be rewritten as y=14-2x subtracting 2x from both sides OR y=-2x+14 clearly you can see that your slope m=-2=-2/1 or 2/-1 So your "rise" is -2 and your "run" is 1 And you can see that your y-intercept is 14 (if you cannot identify this set x=0 and you get y=14) so you have the point (0,14) So if you plot your y-intercept you can get another point on the line by simply going down 2 units and going over to the right 1 unit (using the slope -2/1) And voila, you have two points now and you can draw a line through them so your answer is the step " plot the intercept (7,0) " is not a step when graphing the equation 2x+y=14